# Neural cyberattacks applied to the vision under realistic visual stimuli

### v1

@@ 7 Conclusion
Future work could explore the identification of new cyberattacks focused on other model components, such as the connections between neurons. In the same way, there is a need for works testing the impact and effectiveness of neural cyberattacks on excitatory and inhibitory neurons, over different layers of V1, and distinct neuronal population. In addition, further work is necessary to evaluate the reaction of neurons in the face of color images and static gratings to compare with the movie and drifting grating stimulus used in this paper, respectively. New research also needs to assess other impact metrics to analyze other aspects of neural behavior. Finally, to obtain novel conclusions regarding the impact that these cyberattacks could generate in the real world, it is necessary to keep a close collaboration with neuroscience specialists to evaluate and recreate certain attack situations in real life, such as temporary blindness, or generate stimuli where there were none.

### v2

@@ 7 Conclusion
Future work could explore the identification of new cyberattacks focused on other model components, such as the connections between neurons. Besides, implementing the remaining six neural cyberattacks from the literature would enable a more comprehensive analysis of the effects these types of attacks can have on neural activity. In the same way, there is a need for works testing the impact and effectiveness of neural cyberattacks on excitatory and inhibitory neurons, over different layers of V1, and distinct neuronal population. In addition, further work is necessary to evaluate the reaction of neurons in the face of color images and static gratings to compare with the movie and drifting grating stimulus used in this paper, respectively. New research also needs to assess other impact metrics to analyze other aspects of neural behavior, such as synchronization and group dynamics, including metrics like mean firing rate and population synchrony. Moreover, metrics related to visual discrimination performance, such as reaction time, which measures the time required for neurons to respond to the presentation of a stimulus, would be interesting to use in our study under the effects of the attacks. Finally, to obtain novel conclusions regarding the impact that these cyberattacks could generate in the real world, it is necessary to keep a close collaboration with neuroscience specialists to evaluate and recreate certain attack situations in real life, such as temporary blindness, or generate stimuli where there were none.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
CRediT authorship contribution statement
Victoria Magdalena López Madejska. Methodology, Conceptualization, Investigation, Writing - original draft, Data curation, Software, Visualization. Sergio López Bernal. Methodology, Conceptualization, Writing - Review & Editing. Gregorio Martínez Pérez. Supervision, Project administration, Funding acquisition. Alberto Huertas Celdrán. Methodology, Conceptualization, Funding acquisition, Writing - Review & Editing.